OpenStudy (anonymous):

Multiply and simplify: 2a(7a2 – 3a + 9)

OpenStudy (anonymous):

Alright, if you actually want to do this by hand here's how you can!

OpenStudy (anonymous):

2a(7a^2-3a+9) 1st) 2*7=14-----In regards to a's power, we can note that our exponent rules say to add exponents whenever we multiply, so in this case a+a^2=a^3 2nd) 2*-3=-6-----Same as above a+a=a^2 3rd) 2*9=18-----Same as above a+0=a So combining the steps, we get 14a^3-6a^2+18a

OpenStudy (anonymous):

Thank you soooooooooooooo much. c: That helped a lot! <3